OpenStudy (anonymous):

Evaluate 5x + (-2x 2) for x = 3. a. 21 b. -3 c. 11

OpenStudy (anonymous):

@jacob1233

OpenStudy (whpalmer4):

\[5x-2x^2\]? if \(x=3\) then \[5(3)-2(3)^2 = \]

OpenStudy (anonymous):

i think its -3

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

yes, 15-18 = -3